OpenStudy (anonymous):

Find the axis of symmetry x^2-4x+4=0

OpenStudy (anonymous):

x^2-4x+4=0 =>x^2-2x-2x+4=0 =>x(x-2)-2(x-2)=0 =>(x-2)(x-2)=0 =>x-2=0,x-2=0 x=2 and x=2